One number is four times another number. Determine the two numbers of the sum of their reciprocals is 5/16


1
Expert's answer
2021-08-31T04:30:46-0400

Let's take the two numbers to be "x" and "y"

"x=4y\\\frac{1}{x}+\frac{1}{y}=\frac{5}{16}"

Substitute for "x"

"\frac{1}{4y}+\frac{1}{y}=\frac{5}{16}\\\frac{1+4}{4y}=\frac{5}{16}\\\frac{5}{4y}=\frac{5}{16}\\y=4\\x=4y\\x=4(4)\\x=16"